OpenStudy (anonymous):

Mary deposits $15,000 into a savings account that pays 3.75% interest compounded daily. What will the balance be after five years?

OpenStudy (amistre64):

id just use the compound formula

OpenStudy (amistre64):

\[A=P(1+r/n)^{nt}\]

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

assuming that there are 365 days in a year... then \(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$15,000\\ r=rate\to 3.75\%\to \frac{3.75}{100}\to 0.0375\\ n=\textit{times it compounds per year}\to 365\\ t=years\to 5\)